OpenStudy (erinweeks):

Suppose that a sphere with radius 5a has the same volume as a cone of radius 3a. What is the height of the cone? Give your answer in terms of a. Use the formulas and for the volumes of a sphere and cone respectively.

OpenStudy (erinweeks):

\[V = \frac{ 4 }{ 3 }\pi r ^{3}\]

OpenStudy (erinweeks):

\[V = \frac{ 1 }{ 3 }\pi r ^{2} h\]

OpenStudy (erinweeks):

Helppp .